# Can LLM-based Financial Investing Strategies Outperform the Market in Long Run?

### v2

@@ 6.1 Pitfalls of Selective Evaluation
6.2 Fair Comparisons with the Composite Approach
Selection Strategy Type Timing Strategy SPR $\uparrow$ STR $\uparrow$ AR $\uparrow$ MDD $\uparrow$ AV $\downarrow$
Random 5 S&P 500 (91 symbols) Rule Based Buy and Hold 0.315 0.456 6.694 -35.130 27.410
SMA Cross -0.298 -0.290 0.446 -22.292 15.774
WMA Cross -0.299 -0.305 0.232 -22.754 15.528
ATR Band 0.232 0.425 5.119 -21.535 16.113
Bollinger Bands 0.129 0.288 3.521 -22.487 16.290
Trend Following -0.389 -0.198 2.525 -8.587 8.223
Turn of The Month 0.015 0.072 2.870 -18.582 13.542
Predictor ARIMA 0.255 0.434 6.928 -21.691 17.504
XGBoost -0.055 0.028 3.089 -17.160 13.075
RL A2C 0.086 0.122 1.902 -9.220 6.887
DDPG 0.089 0.125 0.844 -15.185 11.624
PPO 0.179 0.256 3.282 -18.395 13.783
SAC 0.097 0.142 1.389 -16.058 12.375
TD3 0.173 0.248 3.682 -14.471 11.565
LLM FinMem -0.253 0.114 -0.094 -24.243 21.214
FinAgent 0.094 0.323 4.477 -28.059 26.387
Momentum Factor (84 symbols) Rule Based Buy and Hold 0.384 0.694 9.916 -32.596 37.421
SMA Cross -0.251 0.008 2.109 -19.438 20.050
WMA Cross -0.169 0.051 3.674 -18.651 20.330
ATR Band 0.197 0.595 4.314 -19.407 20.038
Bollinger Bands 0.114 0.702 1.881 -19.451 21.555
Trend Following 0.119 0.531 6.380 -15.726 18.696
Turn of The Month 0.056 0.662 3.197 -18.108 18.055
Predictor ARIMA 0.542 1.043 13.257 -18.277 22.892
XGBoost 0.094 1.525 6.131 -12.754 17.238
RL A2C 0.105 0.171 2.488 -14.452 14.815
DDPG 0.209 0.366 5.973 -14.255 16.499
PPO 0.185 0.308 1.939 -23.177 25.527
SAC 0.195 0.321 5.591 -12.235 16.144
TD3 0.186 0.293 3.464 -14.593 14.953
LLM FinMem 0.025 0.170 3.649 -23.335 28.078
FinAgent 0.104 0.534 13.950 -20.675 30.635
Volatility Effect (63 symbols) Rule Based Buy and Hold 0.703 1.291 7.898 -14.146 14.720
SMA Cross -0.568 -0.544 0.781 -9.296 8.665
WMA Cross -0.665 -0.348 1.908 -8.481 8.573
ATR Band -0.026 0.120 2.798 -8.032 7.951
Bollinger Bands -0.077 0.029 2.503 -7.618 7.774
Trend Following 0.230 0.619 5.503 -8.115 9.297
Turn of The Month -0.156 -0.095 2.881 -6.889 7.233
Predictor ARIMA 0.325 0.838 4.898 -9.111 9.807
XGBoost -0.108 -0.055 2.775 -6.676 7.077
RL A2C 0.421 0.795 4.620 -4.428 5.149
DDPG 0.373 0.725 3.768 -6.681 6.693
PPO 0.514 0.972 5.805 -8.757 9.461
SAC 0.402 0.810 3.527 -4.821 5.030
TD3 0.269 0.394 4.610 -5.442 5.992
LLM FinMem -0.228 0.483 4.061 -10.860 11.641
FinAgent 0.241 0.527 4.954 -10.268 11.502
Table 4: Backtest performance under the Composite setup, using three different selection strategies across historical S&P 500 constituents (2004–2024), including delisted symbols. Top in red and second-best in blue.
To overcome the aforementioned biases, we introduce the Composite evaluation setup within FINSABER. This setup integrates systematic selection-based strategies to expand and diversify the stock universe, explicitly addressing survivorship and data-snooping biases. Specifically, we use three unbiased stock selection approaches (details in Appendix B.2): Random Five, Momentum Factor (Muller and and, 2010), and Volatility Effect (Blitz and Vliet, 2007). Each selection strategy is executed at the start of every rolling window, serving as a rebalancing mechanism.
@@ 6.2 Fair Comparisons with the Composite Approach
Table 4 summarises these comprehensive evaluations. Results obtained through this unbiased and systematic approach further validate our previous findings from the selected-four evaluation. Specifically, both the Random Five and Momentum-based selections reinforce the conclusion that the previously claimed superiority of LLM investors is largely driven by selective evaluation setups. For instance, in the Random Five setup, Buy and Hold, ATR Band and ARIMA outperform FinMem and FinAgent in terms of risk-adjusted metrics. Similarly, ARIMA and simple rule-based strategies often perform better than LLM-based methods under the Momentum-based selection. In the Volatility-based selection, traditional methods dominate even more clearly: Buy and Hold achieves the highest Sharpe (0.703), stability (1.291), and AR (7.898%), while PPO and ARIMA again show strong all-round performance. LLM-based methods lag behind, with FinAgent offering moderate returns but lower Sharpe (0.241) and larger drawdowns. Notably, our reported LLM performances do not adjust for potential data leakage: given the use of pretrained models like GPT-4o, the LLMs may have seen parts of the data during training—yet they still fail to outperform traditional strategies under fair evaluation, casting further doubt on their real-world advantage.
@@
Moreover, by comparing Buy and Hold with different selection strategies, we clearly identify the relative effectiveness of each selection strategy: Volatility selection (Sharpe 0.703) outperforms Momentum (0.384), which in turn surpasses Random Five (0.315). RL-based methods exhibit the clearest alignment with selection quality. Strategies like PPO, SAC, and TD3 systematically achieve their best performance under the Volatility selection and degrade under the other two. This suggests RL methods are more sensitive to the quality of the stock candidates. Among LLM strategies, FinAgent exhibits a greater dependency on selection quality than FinMem.
@@
Overall, these results not only confirm our earlier insights but also underscore the critical importance of unbiased, systematic stock-selection methodologies for accurately assessing the true capabilities and limitations of LLM-based investing strategies.
7 Market Regime Analysis
@@
Figure 2: Average Sharpe ratio by regime for all benchmarking strategies. Green = strong, red = weak.
@@
We label each calendar year based on the annual return of the S&P 500: $R_{y}=\frac{P_{T}-P_{0}}{P_{0}}$, where $P_{0}$ and $P_{T}$ are the adjusted closing prices on the first and last trading days of year $y$. A year is classified as bull if $R_{y}\geq+20\%$, bear if $R_{y}\leq-20\%$, and sideways otherwise. The ±$\pm 20\%$ threshold follows standard industry convention (Zweig, 2019).
@@
The LLM strategies perform poorly. FinAgent records Sharpe 0.12 in bulls and -0.38 in bears; FinMem slides from -0.19 to -0.97. Both are too cautious when risk is rewarded and too aggressive when it is penalised. FinAgent is the better of the two, halving the bear-market shortfall relative to Buy and Hold and keeping a small positive Sharpe in neutral conditions, but it still trails every rule-based or predictor benchmark.
@@
8 Findings and Takeaways
The evaluations in §6 and §7 reveals structural limitations in current LLM-based investing strategies. These findings expose methodological flaws in recent studies and highlight the need for more rigorous and domain-aware evaluation.
First, the fragility of LLM-derived alpha under realistic conditions aligns with the Efficient Market Hypothesis (EMH) (Fama, 1970): asset prices reflect all available information, making it difficult to consistently achieve above-average returns through information-based trading. Despite access to large textual corpora, current LLMs fail to extract signals that persist beyond short windows, suggesting that observed gains may stem from narrow evaluation setups rather than genuine alpha.
Second, scaling laws observed in NLP and other domains do not translate effectively to financial applications. While larger models often yield better performance in NLP tasks (Kaplan et al., 2020), our results indicate that increasing model size does not reliably improve investing performance. For instance, FinMem with GPT-4o does not outperform its smaller variant. This suggests that financial markets impose intrinsic limits on extractable signals, regardless of model capacity (Harvey and Liu, 2013). Moreover, simple models like ARIMA or rule-based systems often outperform LLMs on risk-adjusted metrics. This indicates that complexity without financial logic such as trend detection, volatility targeting, or drawdown control, adds little value (and, 2001; Kabiri et al., 2023).
Third, LLMs poorly adapt to market regimes. They tend to be overly conservative in bull markets and excessively aggressive in downturns, contradicting the Adaptive Markets Hypothesis (AMH), which stresses dynamic risk management across shifting conditions (Lo, 2004).
Finally, biased evaluation designs, such as short evaluation periods and selectively chosen symbols, systematically exaggerate results. In finance, this is not merely a technical flaw but a practical risk, as misleading models can drive poor allocation decisions (Sullivan et al., 1999). FINSABER addresses this by enforcing rigorous, bias-aware benchmarking, setting a baseline for credible evaluation in the financial LLM field.
9 Conclusion
@@
We reassess the robustness of LLM timing-based investing strategies using FINSABER, a comprehensive framework that mitigates backtesting biases and extends both the evaluation horizon and symbol universe. Our results show that the perceived superiority of LLM-based methods deteriorates under more robust and broader long-term testing. Regime analysis further reveals that current strategies miss upside in bull markets and incur heavy losses in bear markets due to poor risk control.
@@
We identify two priorities for future LLM-based investors: (1) enhancing uptrend detection to match or exceed passive exposure, and (2) including regime-aware risk controls to dynamically adjust aggression. Addressing these dimensions rather than increasing framework complexity is the key to building practical, reliable strategies.
@@
Finally, our cost analysis (Appendix F) shows that large-scale LLM backtesting is financially intensive. Future work should pursue cost-efficient model designs and incorporate API costs into performance evaluation.
Limitations
There are several limitations to our current study. First, we did not individually tune the traditional rule-based strategies for each rolling evaluation window. Typically, applying domain-specific market insights to optimise parameters can significantly enhance the performance of these methods. However, we argue that our current configuration remains valid and effectively demonstrates the competitive disadvantage faced by LLM strategies. Indeed, tuning the parameters of traditional rule-based strategies would likely elevate their performance further, reinforcing rather than undermining our main conclusions.
Second, our evaluation has not fully eliminated look-ahead bias. Pre-trained LLMs, due to their inherent training corpus, may inadvertently contain stock-related information from historical periods overlapping our test sets. Despite this potential data leakage, the observed underperformance of LLM strategies strengthens our critical assessment. Explicitly addressing this look-ahead concern through controlled model training or careful exclusion of financial data from training corpora will be an important avenue for future research.
Third, to ensure experiment reproducibility, we restricted our analysis to publicly available data, excluding proprietary sources such as private newsfeeds, earning transcripts, or expert analyses. Nonetheless, the FINSABER framework was deliberately designed to be modular and extensible, allowing researchers with access to private data to easily integrate additional information sources. Our primary goal remains providing a rigorous, long-term evaluation pipeline that minimises selective reporting. Researchers lacking proprietary data can fully replicate our results using openly accessible resources.

### v3

@@ 6.1. Pitfalls of Selective Evaluation
6.2. Fair Comparisons with the Composite Approach
To overcome the aforementioned biases, we introduce the Composite evaluation setup within FINSABER. This setup integrates systematic selection-based strategies to expand and diversify the stock universe, explicitly addressing survivorship and data-snooping biases. Specifically, we use four unbiased stock selection approaches from the strategies base (details in Appendix B): Random Five, Momentum Factor (Muller and and, 2010), Volatility Effect (Blitz and Vliet, 2007), and the FinCon Selection Agent in the FinCon (Yu et al., 2024) framework.
For each rolling window, the selection strategy identifies a set of $K$ symbols. Each timing-based strategy is then applied independently to each selected symbol, generating separate trades and performance records. The reported results for each timing strategy reflect the average performance across all selected symbols within the window, as these models operate on individual stocks and do not construct or manage a coordinated portfolio across symbols.
@@ 6.2. Fair Comparisons with the Composite Approach
Type Timing Strategy Random 5 (91 symbols) Momentum Factor (84 symbols)
SPR $\uparrow$ STR $\uparrow$ AR $\uparrow$ MDD $\uparrow$ AV $\downarrow$ SPR $\uparrow$ STR $\uparrow$ AR $\uparrow$ MDD $\uparrow$ AV $\downarrow$
Rule
Based
Buy and Hold 0.315 0.456 6.694 -35.130 27.410 0.384 0.694 9.916 -32.596 37.421
SMA Cross -0.298 -0.290 0.446 -22.292 15.774 -0.251 0.008 2.109 -19.438 20.050
WMA Cross -0.299 -0.305 0.232 -22.754 15.528 -0.169 0.051 3.674 -18.651 20.330
ATR Band 0.232 0.425 5.119 -21.535 16.113 0.197 0.595 4.314 -19.407 20.038
Bollinger Bands 0.129 0.288 3.521 -22.487 16.290 0.114 0.702 1.881 -19.451 21.555
Trend Following -0.389 -0.198 2.525 -8.587 8.223 0.119 0.531 6.380 -15.726 18.696
Turn of The Month 0.015 0.072 2.870 -18.582 13.542 0.056 0.662 3.197 -18.108 18.055
Predictor ARIMA 0.255 0.434 6.928 -21.691 17.504 0.542 1.043 13.257 -18.277 22.892
XGBoost -0.055 0.028 3.089 -17.160 13.075 0.094 1.525 6.131 -12.754 17.238
RL A2C 0.086 0.122 1.902 -9.220 6.887 0.105 0.171 2.488 -14.452 14.815
PPO 0.179 0.256 3.282 -18.395 13.783 0.185 0.308 1.939 -23.177 25.527
SAC 0.097 0.142 1.389 -16.058 12.375 0.195 0.321 5.591 -12.235 16.144
TD3 0.173 0.248 3.682 -14.471 11.565 0.186 0.293 3.464 -14.593 14.953
LLM FinMem -0.253 0.114 -0.094 -24.243 21.214 0.025 0.170 3.649 -23.335 28.078
FinAgent 0.094 0.323 4.477 -28.059 26.387 0.104 0.534 13.950 -20.675 30.635
Type Timing Strategy Volatility Effect (63 symbols) FinCon Selection Agent (80 symbols)
SPR $\uparrow$ STR $\uparrow$ AR $\uparrow$ MDD $\uparrow$ AV $\downarrow$ SPR $\uparrow$ STR $\uparrow$ AR $\uparrow$ MDD $\uparrow$ AV $\downarrow$
Rule
Based
Buy and Hold 0.703 1.291 7.898 -14.146 14.720 0.389 0.671 6.940 -30.943 41.710
SMA Cross -0.568 -0.544 0.781 -9.296 8.665 -0.346 -0.351 -4.187 -21.095 20.765
WMA Cross -0.665 -0.348 1.908 -8.481 8.573 -0.176 -0.129 -1.683 -19.432 21.141
ATR Band -0.026 0.120 2.798 -8.032 7.951 0.181 0.539 4.469 -18.827 24.820
Bollinger Bands -0.077 0.029 2.503 -7.618 7.774 0.116 0.333 7.155 -19.145 27.250
Trend Following 0.230 0.619 5.503 -8.115 9.297 -0.008 0.189 1.358 -19.500 20.400
Turn of The Month -0.156 -0.095 2.881 -6.889 7.233 0.013 0.141 2.020 -15.871 16.862
Predictor ARIMA 0.325 0.838 4.898 -9.111 9.807 0.532 0.841 10.662 -16.018 19.181
XGBoost -0.108 -0.055 2.775 -6.676 7.077 0.116 0.325 8.057 -15.320 18.078
RL A2C 0.421 0.795 4.620 -4.428 5.149 -0.004 -0.061 0.823 -12.557 11.767
PPO 0.514 0.972 5.805 -8.757 9.461 0.132 0.147 2.327 -9.744 10.257
SAC 0.402 0.810 3.527 -4.821 5.030 0.180 0.279 2.661 -11.979 14.210
TD3 0.269 0.394 4.610 -5.442 5.992 0.130 0.334 0.695 -14.621 21.693
LLM FinMem -0.228 0.483 4.061 -10.860 11.641 -0.292 0.135 -1.686 -20.809 24.948
FinAgent 0.241 0.527 4.954 -10.268 11.502 -0.076 0.381 5.168 -15.563 22.565
Table 4. Backtest performance under the Composite setup, using three different selection strategies across historical S&P 500 constituents (2004–2024), including delisted symbols. Top in red and second-best in blue.
Table 4 summarises these comprehensive evaluations. Results obtained through this unbiased and systematic approach further validate our previous findings from the selected-four evaluation. Specifically, both the Random Five and Momentum-based selections reinforce the conclusion that the previously claimed superiority of LLM investors is largely driven by selective evaluation setups. For instance, in the Random Five setup, Buy and Hold, ATR Band and ARIMA outperform FinMem and FinAgent in terms of risk-adjusted metrics. Similarly, ARIMA and simple rule-based strategies often perform better than LLM-based methods under the Momentum-based selection. In the Volatility-based selection, traditional methods dominate even more clearly: Buy and Hold achieves the highest Sharpe (0.703), Sortino (1.291), and AR (7.898%), while PPO and ARIMA again show strong all-round performance. LLM-based methods lag behind, with FinAgent offering moderate returns but lower Sharpe (0.241) and larger drawdowns. Notably, our reported LLM performances do not adjust for potential data leakage: given the use of pretrained models like GPT-4o, the LLMs may have seen parts of the data during training—yet they still fail to outperform traditional strategies under fair evaluation, casting further doubt on their real-world advantage.
@@
Moreover, by comparing Buy and Hold with different selection strategies, we clearly identify the relative effectiveness of each selection strategy: Volatility Effect selection (Sharpe 0.703) outperforms FinCon Selection Agent (0.389) and Momentum Factor (0.384), which in turn surpass Random Five (0.315). RL-based methods exhibit the clearest alignment with selection quality. Strategies like PPO, SAC, and TD3 systematically achieve their best performance under the Volatility selection and degrade under the other three. This suggests RL methods are more dependent on the quality of the stock candidates. Among LLM strategies, FinAgent exhibits a greater dependency on selection quality than FinMem.
@@
Overall, these results not only confirm our earlier insights but also underscore the critical importance of unbiased, systematic stock-selection methodologies for accurately assessing the true capabilities of LLM-based investing strategies.
6.3. Statistical Validation and Behavioural Diagnostics of LLM Agents
To validate our findings from the composite backtests and diagnose the underlying drivers of LLM agent performance, we conduct a unified statistical and behavioural analysis. First, we conduct paired t-tests comparing Buy and Hold, FinMem, and FinAgent across both Selected 4 (Table 3) and Composite (Table 4) setups. Second, we dissect the agents’ behavioural characteristics by examining their drawdown profiles, alpha ($\alpha$) and beta ($\beta$) decomposition, and trading turnover across the different selection environments. These metrics are obtained by regressing the strategy’s excess returns against the market’s excess returns based on the Capital Asset Pricing Model (CAPM) (Sharpe, 1964). The model is defined as: $R_{s}-R_{f}=\alpha+\beta(R_{m}-R_{f})+\epsilon$, where $R_{s}$ is the return of the strategy, $R_{m}$ is the market return, $R_{f}$ is the risk-free rate, and $\epsilon$ is the idiosyncratic residual. In this model, $\beta$ measures the strategy’s systematic risk or volatility relative to the market, while $\alpha$ represents the portion of the return not explained by market exposure, often considered a measure of strategy-specific skill.
Setup B&H vs FinMem B&H vs FinAgent FinMem vs FinAgent
Selective symbols, expanded period (Selected four; Table 3)
TSLA 0.3643 0.1663 0.2258
NFLX 0.0436 0.0363 0.1493
AMZN 0.0127 0.0984 0.4023
MSFT 0.0005 0.2252 0.5549
Bias-mitigated (Composite; Table 4)
Random 5 3.0e-6 7.7e-4 4.0e-3
Momentum 4.0e-5 0.0117 0.2001
Volatility Effect 4.0e-6 5.9e-4 3.8e-3
Table 5. Paired t-test p-values comparing Buy and Hold (B&H), FinMem, and FinAgent under Selected 4 and Composite setups.
Table 5 reports t-tests and p-values for the previous results, testing the null hypothesis of equal performance distributions. Under the selective period, statistical significance is inconsistent and limited mostly to individual stocks. However, after mitigating biases through the composite setup, the p-values drop substantially, indicating the market baseline (Buy and Hold) significantly outperforms both LLM strategies across all robust setups. Notably, while FinAgent tends to outperform FinMem when biases are controlled, both still significantly underperform simple market baselines. Furthermore, the behavioural analysis in Table 6 reveals that this underperformance is rooted in a lack of genuine skill; neither LLM agent generates statistically significant alpha, with all measured p-values exceeding 0.34. This finding robustly supports our main thesis that the claimed superiority of these models does not hold under rigorous evaluation, aligning with the Efficient Market Hypothesis (Lo, 2004).
A clear behavioural hierarchy also emerges between the two agents. FinMem consistently exhibits a more pathological trading profile, characterised by hyperactive turnover and poor risk management. Its commission ratio is approximately 5 to 9 times higher than FinAgent’s across both contexts, and its drawdown durations are substantially longer. This excessive, costly trading correlates with consistent value destruction, as evidenced by FinMem’s negative alpha in all scenarios. In contrast, FinAgent demonstrates a more controlled, albeit still unskilled, approach. Appendix F provides a comparative analysis with visualisations to further highlight the behavioural differences between FinMem and FinAgent.
These behaviours are directly modulated by the selection strategy, which acts as a powerful environmental filter. The Momentum selection strategy elicits the most engaged market posture from the agents, prompting their highest $\beta$ values. FinMem’s performance improves in this context relative to other environments, but it still yields a negative alpha of -1.34%. This is the only scenario where FinAgent produces a large positive alpha of +6.57%. Although this result lacks statistical significance (p=0.35), it suggests that the LLMs’ primary strength may not be in discovering novel signals but rather in exploiting strong, pre-existing market trends. In contrast, the Low Volatility environment prompts a risk-averse posture. Here, FinMem remains ineffective with a -1.04% alpha and a very low $\beta$ of 0.20. FinAgent also becomes highly conservative, with its risk profile improving (e.g., its average drawdown duration falls to 38.71 days) but at the cost of performance, generating a negative alpha.
In summary, this unified analysis statistically validates the underperformance of LLM agents and reveals that their behaviour is not monolithic. It is highly dependent on the characteristics of the asset universe they operate within, reinforcing the need for bias-mitigated evaluation frameworks like FINSABER.
Strategy Avg Max Avg Regular Alpha (%) Beta Alpha
Drawdown (Days) Drawdown (Days) p-value
Momentum Factor
FinMem 210 80 -1.343 0.518 0.477
FinAgent 150 59 6.571 0.758 0.345
Volatility Effect
FinMem 177 71 -1.036 0.199 0.430
FinAgent 123 39 -0.196 0.354 0.368
Table 6. Behavioural analysis of LLM timing strategies, highlighting drawdown duration, alpha ($\alpha$) and beta ($\beta$) decomposition, and trading turnover (commission ratio).
7. Market Regime Analysis
@@
Figure 2. Average Sharpe ratio by regime for all benchmarking strategies. Green = strong, red = weak.
@@
We label each calendar year based on the annual return of the S&P 500: $R_{y}=\frac{P_{T}-P_{0}}{P_{0}}$, where $P_{0}$ and $P_{T}$ are the adjusted closing prices on the first and last trading days of year $y$. A year is classified as bull if $R_{y}\geq+20\%$, bear if $R_{y}\leq-20\%$, and sideways otherwise. The $\pm 20\%$ threshold follows standard industry convention (Zweig, 2019).
@@
LLM strategies perform poorly. FinAgent records Sharpe 0.12 in bulls and -0.38 in bears; FinMem gets -0.19 and -0.97. Both are too cautious when risk is rewarded and too aggressive when it is penalised. FinAgent is better, halving the bear-market shortfall relative to Buy and Hold and keeping a small positive Sharpe in neutral conditions, but it still trails rule-based or predictor benchmark.
@@
8. Findings and Takeaways
Our investigation via the FINSABER framework offers several novel findings that challenge the prevailing narrative on LLM-based investors and set a new baseline for future research.
First, we find that LLM-derived alpha is likely a methodological artefact of narrow, biased evaluations. The performance advantages reported in short-term, selective studies vanish under our bias-mitigated backtests, which reveal a consistent and statistically significant failure to generate alpha (Section 6.3). This suggests that current LLMs do not overcome the Efficient Market Hypothesis (Fama, 1970) in realistic conditions, and that prior gains stemmed from survivorship and look-ahead biases rather than genuine market inefficiency.
Second, model complexity does not equate to market competence. The scaling laws of natural language processing (Kaplan et al., 2020) do not translate effectively to financial markets, which impose intrinsic limits on extractable signals (Harvey and Liu, 2013). We show that larger models do not reliably outperform smaller ones, and both are consistently bettered by simpler models like ARIMA on risk-adjusted metrics (Table 4). Without encoded financial logic, architectural complexity appears to add noise rather than value.
Third, we diagnose “how” LLM agents fail, revealing a fundamental misalignment with market regimes. Our analysis (§7, Appendix F) shows that agents are pathologically miscalibrated: they are too conservative in bull markets and too aggressive in bear markets. This behavioural flaw contradicts the Adaptive Markets Hypothesis (Lo, 2004), shifting the issue from merely a lack of profitability to a more profound failure in the agents’ decision-making policies.
Synthesising these points, our work establishes that the primary barrier to successful LLM investors is not model scale, but a lack of domain-aware financial logic. The path forward is designing smarter, more adaptive agents, and FINSABER provides the framework to rigorously test such designs, moving the field beyond flawed evaluations toward practical and robust financial agents.
9. Conclusion
@@
We reassess the robustness of LLM timing-based investing strategies using FINSABER, a comprehensive framework that mitigates backtesting biases and extends both the evaluation horizon and symbol universe. Results show that the perceived superiority of LLM-based methods deteriorates under more robust and broader long-term testing. Regime analysis further reveals that current strategies miss upside in bull markets and incur heavy losses in bear markets due to poor risk control.
@@
We identify two priorities for future LLM-based investors: (1) enhancing uptrend detection to match passive exposure, and (2) including regime-aware risk controls to dynamically adjust aggression. Addressing these dimensions rather than increasing framework complexity is the key to building practical, reliable strategies.
@@
Finally, our cost analysis (Appendix G) shows that large-scale LLM backtesting is financially intensive. Future work should pursue cost-efficient model designs and incorporate API costs into performance evaluation.